OpenStudy (anonymous):

5x^3+19x^2+16x-10 / 5x-1. Find the quotient and the remainder using long divison. If there is a remainder type your answer in the form quotient + remainder / divisor.

jimthompson5910 (jim_thompson5910):

Use polynomial long division to get (see attached) Quotient: \(\Large x^2+4x+4 \) Remainder: \(\Large -6\) So \[\Large \frac{5x^3+19x^2+16x-10}{5x-1} = x^2+4x+4 + \frac{-6}{5x-1}\]

OpenStudy (anonymous):

so that's the answer?

jimthompson5910 (jim_thompson5910):

yes